OpenStudy (anonymous):

Medal!!!!!!!!!!! Historical Feature: In the equation x^3+px+q=o, replace x by H+K. Let 3HK=-p, and show that H^3+K^3=-q

OpenStudy (perl):

$$ \Large{ x^3+px+q=0 \\ x= h+k \\ 3hk=-p \\~\\ \textrm{Our goal is show that} ~h^3+k^3=-q \\ \textrm{Steps: substitute} \\~\\ (h+k)^3 + (-3hk)(h+k) + q=0 } $$

OpenStudy (perl):

can you expand that?

OpenStudy (anonymous):

q=-h^3-k^3

OpenStudy (perl):

thats what you need to show

OpenStudy (anonymous):

wait what